[Date Prev][Date Next][Thread Prev][Thread Next] - [Date Index][Thread Index][Author Index]

OSCAR-11 Report

                OSCAR-11 REPORT    15 April 1998

In response  to  many  requests  for  information  about  methods  of
decoding  OSCAR-11 signals, I have now added a package of information
to my web  site.  The  web  site  also  contains  some  software  for
capturing  data,  and  decoding  ASCII telemetry and WOD. There is an
archive of raw data (mainly WOD) for analysis, which  is  continually
being expanded, as new data is captured.

There  are  also  some  audio  files, including a recording of mode-S
which plays for 20 seconds.  The other audio files  are  examples  of
each  type  of  data  transmitted by OSCAR-11, and each one plays for
about ten seconds. All the audio files are zipped, so that  they  can
be  played off-line.  They should help listeners identify the various
types of data, and give an indication of the signal quality  required
for successful decoding. The URL is -


When  the  previous  OSCAR-11  report  was  written,  ground  control
operations  were  taking place to investigate a 'shut down' which had
occurred a few days previously.  These operations were successful and
normal service has been restored. During the period 21  March  to  15
April  excellent  signals  have  once  again  been  received from the
145.826 MHz. beacon.

Three  reports  of  the  mode-S  beacon have been received, from Mike
WL7BQQM, Ed K9EK, and Jim KK3K.  These reports  were  all  after  the
'shut  down'  had occurred, indicating that beacon is still OK.  Many
thanks for those reports.

Since  the  'shut  down'  the  telemetry  indicates  that  there is a
reduction in the 14 volt line current of about 50 mA and a  reduction
in  the +5 volt current of about 25 mA.  The reason for these changes
is not obvious (to me), as the status telemetry does  not  show  that
any  units have been switched OFF.  The only obvious change in status
telemetry  is  that  the  four  bits  associated  monitoring   camera
integration  period and gain have changed from ONEs to ZEROs. However
the camera power is OFF (unchanged).

For  several years I have been monitoring the efficiency of the power
conversion process, ie the ratio of power output to power input.

Power input = Sum of solar array currents * Array voltage

Power output = Battery volts * (Charge current + 14volt line current)

The efficiency calculated in this way has usually been around  80  to
85%. However since the 'shut down' it has fallen to around 65 to 70%.
I  would  welcome  any  explanations  or  suggestions  regarding this

A new WOD survey of  channels  1,  2,  3,  61  (magnetometers)  dated
19-March-98  has  been started.  Anyone using this survey should note
the unusual starting time of 16:00:05 UTC.

The operating schedule is unchanged.

        ASCII status (210 seconds)
        ASCII bulletin  (60 seconds)
        BINARY SEU (30 seconds)
        ASCII TLM (90 seconds)
        ASCII WOD (120 seconds)
        ASCII bulletin (60 seconds)
        BINARY ENG (30 seconds)

There  are  also  additional  status  blocks  after  each bulletin is
transmitted, and between ASCII TLM and WOD.

The mode-S beacon is ON, transmitting  an  unmodulated  carrier,  but
telemetry indicates that it has partially failed, and delivering half
power.   This beacon is a useful test source for those testing mode-S
converters, prior to the launch of P3-D.  It is  considerably  weaker
than  DOVE,  which should be used for initial testing. Any reports of
reception  on  2401  MHz.   would  be  most  welcome.  Please  e-mail

The 435.025 MHz. beacon is normally OFF.  However it can sometimes be
heard  when  the  satellite is being commanded by ground control, ie.
within range of Guildford, UK.  When the 435 beacon is  transmitting,
the  145  beacon  is  normally  OFF.   The data transmitted is mainly

If you place this bulletin on a terrestrial  packet  network,  please
use the bulletin identifier $BID:U2RPT24.CWV, to prevent duplication.

73 Clive G3CWV   g3cwv@amsat.org